OpenStudy (anonymous):

The sum of 43.9 and a number is 49.65, as shown below. What number should go in the box to complete the addition problem? 43.9 plus box.75 equals 49.65

OpenStudy (anonymous):

49.65 - 43.9 = [ ] .75 if you subtract them out, you get 5.75, so put 5 in the box

OpenStudy (anonymous):

Solve the following for x:\[x+\frac{439}{10}+\frac{75}{100}=\frac{4965}{100} \]x = 5